# Umlaut information

### v1

@@ IV.5 Operational interpretation in unassisted communication
$E^{\emptyset}(0^{+},\pazocal{W})=-\min_{P_{X}}\sum_{x,x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}.$ (121)
@@
$E^{\emptyset}(0^{+},\pazocal{W})\leq E^{\rm NS}(0^{+},\pazocal{W}).$ (122)
@@
$-\sum_{x,x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}\leq-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right).$ (123)
@@
$-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)$ $=\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})$ (124)
@@
$\min_{Q_{Y}}\sum_{x\in\mathcal{X}}$ $P_{X}(x)D\big{(}Q_{Y}\|\pazocal{W}(\,\cdot\,|x)\big{)}$ (125)
@@
$\min_{Q_{Y}}\Bigg{(}\frac{1}{2}$ $D\big{(}Q_{Y}\|\pazocal{W}(\,\cdot\,|x)\big{)}+\frac{1}{2}D\big{(}Q_{Y}\|\pazocal{W}(\,\cdot\,|x^{\prime})\big{)}\Bigg{)}$ (126)
@@
We extend the notion of channel coding to list decoding in order to derive the second operational interpretation of the umlaut information. Given a set of messages =$\mathcal{M}=\{1,\dots,M\}$ and a discrete memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, an $L$-list unassisted code $\pazocal{Z}$ is given by the composition of an encoder $\pazocal{E}$ from $\mathcal{M}$ to $\mathcal{X}$ with a decoder $\pazocal{D}$ from $\mathcal{Y}$ to $[\mathcal{M}]^{L}$, defined as the family of subsets of $\mathcal{M}$ with cardinality $L$. Calling $W$ the source of messages, i.e. a random variable taking values in $\mathcal{M}$, which is assumed to be uniform, let $\hat{W}_{L}$ be the random output of the code taking values in $[\mathcal{M}]^{L}$. Then the transmission is considered to be successful if $W\in\hat{W}_{L}$.
@@
$\varepsilon_{L}^{\emptyset}(M,\pazocal{W})\coloneqq\min_{\pazocal{Z}\in\{L\text{-list unassisted codes}\}}\left\{\mathbb{P}\left[W\neq\hat{W}_{L}\right]\quad\text{ with }\quad Y|X\sim\pazocal{W},\quad X\hat{W}_{L}|YW\sim\pazocal{Z}\right\}.$ (127)
@@
$E_{L}^{\emptyset}(r,\pazocal{W})\coloneqq\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon_{L}^{\emptyset}(L\exp(rn),\pazocal{W})\,.$ (128)
@@
$E^{\emptyset}_{L}(0^{+},\pazocal{W})$ $\coloneqq\lim_{r\to 0^{+}}E_{L}^{\emptyset}(r,\pazocal{W}).$ (129)
@@
$E^{\emptyset}_{L}(0^{+},\pazocal{W})=\max_{P_{X}}\sum_{x_{1},\dots,x_{L+1}}P_{X}(x_{1})\cdots P_{X}(x_{L+1})\left(-\log\sum_{y\in\mathcal{Y}}\sqrt[L+1]{\pazocal{W}(y|x_{1})\cdots\pazocal{W}(y|x_{L+1})}\right),$ (130)
@@
Given a discrete, memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, the umlaut information of $\pazocal{W}$ is an upper bound to its $L$-list unassisted zero-rate error exponent $E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq U(\pazocal{W}),$ (131) with equality in the large list limit. $\sup_{L\geq 1}E^{\emptyset}_{L}(0^{+},\pazocal{W})=U(\pazocal{W}).$ (132)
@@
$\ell_{k,q}(\pazocal{W},P_{X})\coloneqq\sum_{x_{1},\dots,x_{k}\in\mathcal{X}}P_{X}(x_{1})\cdots P_{X}(x_{k})\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{k}q_{i}\log\pazocal{W}(y|x_{i})\right)\right).$ (133)
@@
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\sum_{x_{1},\dots,x_{k}\in\mathcal{X}}P_{X}(x_{1})\cdots P_{X}(x_{k})\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{k}q_{i}\log\pazocal{W}(y|x_{i})\right)\right),$
@@
$U(X;Y)=\sup_{k\geq 1}\ell_{k,u_{k}}(\pazocal{W},P_{X})=\lim_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X}).$ (138)
By inspection, Propositions 31 and 32 taken together immediately imply the sought-after Theorem 29.
@@
$\ell_{L+1,u_{L+1}}(P_{X},\pazocal{W})$ $=\sum_{x_{1},\dots,x_{L+1}\in\mathcal{X}}P_{X}(x_{1})\cdots P_{X}(x_{L+1})\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{L+1}\frac{1}{L+1}\log\pazocal{W}(y|x_{i})\right)\right)$ (139)
$=\sum_{x_{1},\dots,x_{L+1}}P_{X}(x_{1})\cdots P_{X}(x_{L+1})\left(-\log\sum_{y\in\mathcal{Y}}\sqrt[L+1]{\pazocal{W}(y|x_{1})\cdots\pazocal{W}(y|x_{L+1})}\right).$
Taking the maximum over $P_{X}\in\mathcal{P}(\mathcal{X})$ and by Propositions 28 and 31, we have
$E^{\emptyset}_{L}(0^{+},\pazocal{W})=\max_{P_{X}}\ell_{L+1,u_{L+1}}(P_{X},\pazocal{W})\leq U(\pazocal{W}).$ (140)
@@
Let us consider a fixed a discrete memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$. Then the map $f:[0,1]^{\mathcal{X}}\to\mathbb{R}$, defined in terms of Convention 18 as
$f(v)\coloneqq-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}v(x)\log\pazocal{W}(y|x)\right),$ (141)
is monotone increasing, in the sense that, if $v(x)\geq u(x)$ for all $x\in\mathcal{X}$ , then $f(v)\geq f(u)$. Furthermore, it is lower semi-continuous and uniformly lower bounded by
$f\geq-\log|\mathcal{Y}|.$ (142)
@@
According to Convention 18, $f$ is meant to be defined as
$f(v)=-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}v(x)\log\pazocal{W}(y|x)\right),$ (143)
where ${\rm supp}(v)\coloneqq\{x\in\mathcal{X}\,:\,v(x)>0\}$ and $\mathcal{Y}_{\pazocal{W},v}\coloneqq\{y\in\mathcal{Y}\,:\,\pazocal{W}(y|x)>0,\;\forall\,x\in{\rm supp}(v)\}$. Let us suppose that $v(x)\geq u(x)$ for all $x\in\mathcal{X}$. Then, since $\log\pazocal{W}(y|x)\leq 0$ for any $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ and ${\rm supp}(v)\supseteq{\rm supp}(u)$, we have
$\sum_{x\in{\rm supp}(v)}v(x)\log\pazocal{W}(y|x)\leq\sum_{x\in{\rm supp}(u)}u(x)\log\pazocal{W}(y|x)$ (144)
for any $y\in\mathcal{Y}_{\pazocal{W},v}\cap\mathcal{Y}_{\pazocal{W},u}$. Using again that ${\rm supp}(v)\supseteq{\rm supp}(u)$, we see that $\mathcal{Y}_{\pazocal{W},v}\subseteq\mathcal{Y}_{\pazocal{W},u}$. By increasing monotonicity and positivity of the exponential, we get
$\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}v(x)\log\pazocal{W}(y|x)\right)$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{\leq}}\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}u(x)\log\pazocal{W}(y|x)\right)$ (145)
@@
where in (i) we have used (144) and the fact that $\mathcal{Y}_{\pazocal{W},v}\cap\mathcal{Y}_{\pazocal{W},u}=\mathcal{Y}_{\pazocal{W},v}$; in (ii) we have noticed that extending the sum over $\mathcal{Y}_{\pazocal{W},v}$ to a sum over $\mathcal{Y}_{\pazocal{W},u}$ introduces further positive terms. By decreasing monotonicity of $-\log(\,\cdot\,)$ we conclude that $f(v)\geq f(u)$. Leveraging the monotonicity, we conclude that $f(v)\geq f(v_{0})$ where $v_{0}(x)\coloneqq 0$ for any $x\in\mathcal{X}$, hence
$f(v)\geq-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},v_{0}}}1\geq-\log|\mathcal{Y}|.$ (146)
The remaining part to be proved is the lower semicontinuity. For any fixed nonempty subset $\mathcal{X}^{\ast}\subseteq\mathcal{X}$ let us define $V_{\mathcal{X}^{\ast}}\coloneqq\{v\in[0,1]^{\mathcal{X}}\,:\,{\rm supp}(v)=\mathcal{X}^{\ast}\}$. Then we claim that the restriction $f|_{V_{\mathcal{X}^{\ast}}}:V_{\mathcal{X}^{\ast}}\to\mathbb{R}$ is continuous. Indeed, let us first rewrite
$f|_{V_{\mathcal{X}^{\ast}}}(v)=-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},\mathcal{X}^{\ast}}}\exp\left(\sum_{x\in\mathcal{X}^{\ast}}v(x)\log\pazocal{W}(y|x)\right),$ (147)
where we have noticed that the support of $v\in(0,1]^{\mathcal{X}^{\ast}}$ is always $\mathcal{X}^{\ast}$ and we have introduced $\mathcal{Y}_{\pazocal{W},\mathcal{X}^{\ast}}\coloneqq\{y\in\mathcal{Y}\,:\,\pazocal{W}(y|x)>0,\;\forall\,x\in\mathcal{X}^{\ast}\}$. By noticing that $\mathcal{X}$ and $\mathcal{Y}$ are finite, $\mathcal{Y}_{\pazocal{W},\mathcal{X}^{\ast}}$ does not depend on $v\in V_{\mathcal{X}^{\ast}}$, and $f|_{V_{\mathcal{X}^{\ast}}}(v)$ is written as a composition of continuous functions, we conclude that $v\mapsto f|_{V_{\mathcal{X}^{\ast}}}(v)$ is a continuous function. Let us consider now a generic sequence $\{v_{k}\}_{k\in\mathbb{N}}$ in $[0,1]^{\mathcal{X}}$ that converges to $v_{\infty}\in[0,1]^{\mathcal{X}}$. Let $\{u_{k}\}_{k\in\mathbb{N}}$ be the sequence defined as
$u_{k}(x)=\begin{cases}v_{k}(x)&x\in{\rm supp}(v_{\infty})\\ 0&x\notin{\rm supp}(v_{\infty})\end{cases}$ (148)
@@
$f(v_{k})\geq f(u_{k})=f|_{V_{{\rm supp}(v_{\infty})}}(u_{k})\qquad\forall\,k\geq k^{\ast}.$ (149)
@@
$\liminf_{k\to\infty}f(v_{k})\geq\liminf_{k\to\infty}f|_{V_{{\rm supp}(v_{\infty})}}(u_{k})=f|_{V_{{\rm supp}(v_{\infty})}}(v_{\infty})=f(v_{\infty}),$ (150)
@@
$\ell_{k,u_{k}}(\pazocal{W},P_{X})=\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{k}\frac{1}{k}\log\pazocal{W}(y|X_{i})\right)\right]$ (151)
@@
$\sum_{i=1}^{k}\frac{1}{k}\log\pazocal{W}(y|X_{i})=\sum_{x\in\mathcal{X}}\frac{N(x|X^{k})}{k}\log\pazocal{W}(y|x)$ (152)
@@
$\hat{N}_{k}^{x}\coloneqq\frac{N(x|X^{k})}{k}=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{X_{i}=x}$ (153)
which, by the weak law of the large numbers, converges in probability to its expectation value $\mathbb{E}\left[\hat{N}_{k}^{x}\right]=P_{X}(x)$. More precisely, let us fix $0<\varepsilon<\min\{P_{X}(x)\,:\,x\in{\rm supp}(P_{X})\}$, and let us call $\mathcal{E}_{k}^{(\varepsilon)}$ the event $\left\{\exists\,x\in\mathcal{X}:\big{|}\hat{N}_{k}^{x}-P_{X}(x)\big{|}>\varepsilon\right\}$ and $\big{(}\mathcal{E}_{k}^{(\varepsilon)}\big{)}^{c}$ its complementary. Then
$\lim_{k\to\infty}\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)=0$ (154)
@@
$\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $=\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\right]$ (155)
@@
$\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $=\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\right]$ (156)
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\,\middle|\,\mathcal{E}_{k}^{(\varepsilon)}\,\right]\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)$

### v2

@@ IV.6 Operational interpretation in unassisted communication
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})=-\min_{P_{X}}\sum_{x,x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}.$ (161)
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})\leq E^{\rm NS}(0^{+},\pazocal{W}).$ (162)
@@
$\hskip 0.0pt-\sum_{x,x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}\leq-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right).$ (163)
@@
$\hskip 0.0pt-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)$ $=\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})$ (164)
@@
$\hskip 0.0pt\min_{Q_{Y}}\sum_{x\in\mathcal{X}}$ $P_{X}(x)D\big{(}Q_{Y}\|\pazocal{W}(\,\cdot\,|x)\big{)}$ (165)
@@
$\hskip 0.0pt\min_{Q_{Y}}\Bigg{(}\frac{1}{2}$ $D\big{(}Q_{Y}\|\pazocal{W}(\,\cdot\,|x)\big{)}+\frac{1}{2}D\big{(}Q_{Y}\|\pazocal{W}(\,\cdot\,|x^{\prime})\big{)}\Bigg{)}$ (166)
@@
We extend the notion of channel coding to list decoding in order to derive the second operational interpretation of the umlaut information. Given a set of messages $\mathcal{M}=\{1,\dots,M\}$ and a discrete memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, an $L$-list unassisted code $\pazocal{Z}$ is given by the composition of an encoder $\pazocal{E}$ from $\mathcal{M}$ to $\mathcal{X}$ with a decoder $\pazocal{D}$ from $\mathcal{Y}$ to $\mathcal{M}]^{L}$, defined as the family of subsets of $\mathcal{M}$ with cardinality $L$. Calling $W$ the source of messages, i.e. a random variable taking values in $\mathcal{M}$, which is assumed to be uniform, let $\hat{W}_{L}$ be the random output of the code taking values in $\mathcal{M}]^{L}$. Then the transmission is considered to be successful if $W\in\hat{W}_{L}$.
@@
$\hskip 0.0pt\varepsilon_{L}^{\emptyset}(M,\pazocal{W})\coloneqq\min_{\pazocal{Z}\in\{L\text{-list unassisted codes}\}}\left\{\mathbb{P}\left[W\notin\hat{W}_{L}\right]\quad\text{ with }\quad Y|X\sim\pazocal{W},\quad X\hat{W}_{L}|YW\sim\pazocal{Z}\right\}.$ (167)
@@
$\hskip 0.0ptE_{L}^{\emptyset}(r,\pazocal{W})\coloneqq\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon_{L}^{\emptyset}(L\exp(rn),\pazocal{W})\,.$ (168)
@@
$\hskip 0.0ptE^{\emptyset}_{L}(0^{+},\pazocal{W})$ $\coloneqq\lim_{r\to 0^{+}}E_{L}^{\emptyset}(r,\pazocal{W}).$ (169)
@@
[…]
@@
Given a discrete, memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, the umlaut information of $\pazocal{W}$ is an upper bound to its $L$-list unassisted zero-rate error exponent $\hskip 0.0ptE^{\emptyset}_{L}(0^{+},\pazocal{W})\leq U(\pazocal{W}),$ (171) with equality in the large list limit: $\hskip 0.0pt\sup_{L\geq 1}E^{\emptyset}_{L}(0^{+},\pazocal{W})=U(\pazocal{W}).$ (172)
We remark that in our work, the size of the list is fixed when the asymptotic limit of a large number of uses of the channel is taken. In the literature, another setting was also considered: $L$ could be taken dependent on the block length $n$ as $L=\exp(\lambda n)$. This case is considered for instance in [11, Theorem 3], where it is proved that
$\hskip 0.0ptE_{L}^{\emptyset}(r,\pazocal{W})=E_{\rm sp}(r-\lambda,\pazocal{W}).$ (173)
This result can be connected to the umlaut information of $\pazocal{W}$ as the error exponent converges to $U(\pazocal{W})$ if we consider the limit $r\to 0^{+}$ with the constraint $\lambda=\lambda_{r}<r$:
$\hskip 0.0ptE_{L}^{\emptyset}(0^{+},\pazocal{W})=\lim_{r\to 0^{+}}E_{\rm sp}(r-\lambda_{r},\pazocal{W})=E_{\rm sp}(0^{+},\pazocal{W})=U(\pazocal{W}).$ (174)
In our discussion, however, we will focus only on the fixed list size regime. Random coding bounds on the list decoding error exponent in this setting have also been considered in [34], from which connections with the sphere packing bound in the large list limit can also be deduced.
@@
$\hskip 0.0pt\ell_{k,q}(\pazocal{W},P_{X})\coloneqq\sum_{x_{1},\dots,x_{k}\in\mathcal{X}}P_{X}(x_{1})\cdots P_{X}(x_{k})\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{k}q_{i}\log\pazocal{W}(y|x_{i})\right)\right).$ (175)
@@
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\sum_{x_{1},\dots,x_{k}\in\mathcal{X}}P_{X}(x_{1})\cdots P_{X}(x_{k})\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{k}q_{i}\log\pazocal{W}(y|x_{i})\right)\right),$
@@
$\hskip 0.0ptU(X;Y)=\sup_{k\geq 1}\ell_{k,u_{k}}(\pazocal{W},P_{X})=\lim_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X}).$ (180)
By inspection, Propositions 36 and 37 taken together immediately imply the sought-after Theorem 34.
@@
$\hskip 0.0pt\ell_{L+1,u_{L+1}}(P_{X},\pazocal{W})$ $=\sum_{x_{1},\dots,x_{L+1}\in\mathcal{X}}P_{X}(x_{1})\cdots P_{X}(x_{L+1})\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{L+1}\frac{1}{L+1}\log\pazocal{W}(y|x_{i})\right)\right)$ (181)
[…]
Taking the maximum over $P_{X}\in\mathcal{P}(\mathcal{X})$ and by Propositions 33 and 36, we have
$\hskip 0.0ptE^{\emptyset}_{L}(0^{+},\pazocal{W})=\max_{P_{X}}\ell_{L+1,u_{L+1}}(P_{X},\pazocal{W})\leq U(\pazocal{W}).$ (182)
@@
Let us consider a fixed a discrete memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$. Then the map $f:[0,1]^{\mathcal{X}}\to\mathbb{R}\cup\{+\infty\}$, defined in terms of Convention 17 as
$\hskip 0.0ptf(v)\coloneqq-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}v(x)\log\pazocal{W}(y|x)\right),$ (183)
is monotone increasing, in the sense that, if $v(x)\geq u(x)$ for all $x\in\mathcal{X}$ , then $f(v)\geq f(u)$. Furthermore, it is lower semi-continuous and, if $\|v\|_{1}=1$, non-negative:
$\hskip 0.0ptf(v)\geq 0\,.$ (184)
@@
According to Convention 17, $f$ is meant to be defined as
$\hskip 0.0ptf(v)=-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}v(x)\log\pazocal{W}(y|x)\right),$ (185)
where ${\rm supp}(v)\coloneqq\{x\in\mathcal{X}\,:\,v(x)>0\}$ and $\mathcal{Y}_{\pazocal{W},v}\coloneqq\{y\in\mathcal{Y}\,:\,\pazocal{W}(y|x)>0\;\forall\,x\in{\rm supp}(v)\}$. Let us suppose that $v(x)\geq u(x)$ for all $x\in\mathcal{X}$. Then, since $\log\pazocal{W}(y|x)\leq 0$ for any $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ and ${\rm supp}(v)\supseteq{\rm supp}(u)$, we have
$\hskip 0.0pt\sum_{x\in{\rm supp}(v)}v(x)\log\pazocal{W}(y|x)\leq\sum_{x\in{\rm supp}(u)}u(x)\log\pazocal{W}(y|x)$ (186)
for any $y\in\mathcal{Y}_{\pazocal{W},v}\cap\mathcal{Y}_{\pazocal{W},u}$. Using again that ${\rm supp}(v)\supseteq{\rm supp}(u)$, we see that $\mathcal{Y}_{\pazocal{W},v}\subseteq\mathcal{Y}_{\pazocal{W},u}$. By the monotonicity and positivity of the exponential, we get
$\hskip 0.0pt\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}v(x)\log\pazocal{W}(y|x)\right)$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{\leq}}\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}u(x)\log\pazocal{W}(y|x)\right)$ (187)
@@
where in (i) we have used (186); in (ii) we have noticed that extending the sum over $\mathcal{Y}_{\pazocal{W},v}$ to a sum over $\mathcal{Y}_{\pazocal{W},u}$ introduces further positive terms. By anti-monotonicity of $\log(\,\cdot\,)$ we conclude that $f(v)\geq f(u)$. Let $v\in[0,1]^{\mathcal{X}}$ such that $\|v\|_{1}=1$. Then
$\hskip 0.0ptf(v)=-H(v)-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},v}}\exp\left(\sum_{x\in{\rm supp}(v)}v(x)\log\left(v(x)\pazocal{W}(y|x)\right)\right)=\min_{Q_{Y}}D(vQ_{y}\|P^{v}_{\pazocal{W}}),$ (188)
where we have leveraged the Gibbs variational principle (Lemma 1) in the last identity to introduce the minimization over $Q_{Y}\in\mathcal{P}(\mathcal{Y})$ and we have defined $P^{v}_{\pazocal{W}}(x,y)\coloneqq v(x)\pazocal{W}(y|x)$. Then (184) follows from the positivity of the relative entropy between probability distributions. The remaining part to be proved is the lower semicontinuity. For any fixed nonempty subset $\mathcal{X}^{\ast}\subseteq\mathcal{X}$ let us define $V_{\mathcal{X}^{\ast}}\coloneqq\{v\in[0,1]^{\mathcal{X}}\,:\,{\rm supp}(v)=\mathcal{X}^{\ast}\}$. Then we claim that the restriction $f|_{V_{\mathcal{X}^{\ast}}}:V_{\mathcal{X}^{\ast}}\to\mathbb{R}$ is continuous. Indeed, let us first rewrite
$\hskip 0.0ptf|_{V_{\mathcal{X}^{\ast}}}(v)=-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},\mathcal{X}^{\ast}}}\exp\left(\sum_{x\in\mathcal{X}^{\ast}}v(x)\log\pazocal{W}(y|x)\right),$ (189)
where we have noticed that the support of $v\in(0,1]^{\mathcal{X}^{\ast}}$ is always $\mathcal{X}^{\ast}$ and we have introduced $\mathcal{Y}_{\pazocal{W},\mathcal{X}^{\ast}}\coloneqq\{y\in\mathcal{Y}\,:\,\pazocal{W}(y|x)>0,\;\forall\,x\in\mathcal{X}^{\ast}\}$. By noticing that $\mathcal{X}$ and $\mathcal{Y}$ are finite, $\mathcal{Y}_{\pazocal{W},\mathcal{X}^{\ast}}$ does not depend on $v\in V_{\mathcal{X}^{\ast}}$, and $f|_{V_{\mathcal{X}^{\ast}}}(v)$ is written as a composition of continuous functions, we conclude that $v\mapsto f|_{V_{\mathcal{X}^{\ast}}}(v)$ is a continuous function. Let us consider now a generic sequence $\{v_{k}\}_{k\in\mathbb{N}}$ in $0,1]^{\mathcal{X}}$ that converges to $v_{\infty}\in[0,1]^{\mathcal{X}}$. Let $\{u_{k}\}_{k\in\mathbb{N}}$ be the sequence defined as
[…]
@@
$\hskip 0.0ptf(v_{k})\geq f(u_{k})=f|_{V_{{\rm supp}(v_{\infty})}}(u_{k})\qquad\forall\,k\geq k^{\ast}.$ (191)
@@
$\hskip 0.0pt\liminf_{k\to\infty}f(v_{k})\geq\liminf_{k\to\infty}f|_{V_{{\rm supp}(v_{\infty})}}(u_{k})=f|_{V_{{\rm supp}(v_{\infty})}}(v_{\infty})=f(v_{\infty}),$ (192)
@@
$\hskip 0.0pt\ell_{k,u_{k}}(\pazocal{W},P_{X})=\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{i=1}^{k}\frac{1}{k}\log\pazocal{W}(y|X_{i})\right)\right]$ (193)
@@
$\hskip 0.0pt\sum_{i=1}^{k}\frac{1}{k}\log\pazocal{W}(y|X_{i})=\sum_{x\in\mathcal{X}}\frac{N(x|X^{k})}{k}\log\pazocal{W}(y|x)$ (194)
@@
$\hskip 0.0pt\hat{N}_{k}^{x}\coloneqq\frac{N(x|X^{k})}{k}=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{X_{i}=x}$ (195)
which, by the weak law of the large numbers, converges in probability to its expectation value $\mathbb{E}\left[\hat{N}_{k}^{x}\right]=P_{X}(x)$. More precisely, let us fix $0<\varepsilon<\min\{P_{X}(x)\,:\,x\in{\rm supp}(P_{X})\}$, and let us call $\mathcal{E}_{k}^{(\varepsilon)}$ the event $\left\{\exists\,x\in\mathcal{X}:\big{|}\hat{N}_{k}^{x}-P_{X}(x)\big{|}>\varepsilon\right\}$ and $\big{(}\mathcal{E}_{k}^{(\varepsilon)}\big{)}^{c}$ its complement. Then
$\hskip 0.0pt\lim_{k\to\infty}\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)=0$ (196)
@@
$\hskip 0.0pt\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $=\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\right]$ (197)
@@
$\hskip 0.0pt\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\,\middle|\,\mathcal{E}_{k}^{(\varepsilon)}\,\right]\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)$ (198)